hOW TO GET SUM 50 OF ODD 1 TO 19 ?
Answer by Edwin McCravy(20086)

1 + 13 + 17 + 19 = 50
3 + 11 + 17 + 19 = 50
3 + 13 + 15 + 19 = 50
5 + 9 + 17 + 19 = 50
5 + 11 + 15 + 19 = 50
5 + 13 + 15 + 17 = 50
7 + 9 + 15 + 19 = 50
7 + 11 + 13 + 19 = 50
7 + 11 + 15 + 17 = 50
9 + 11 + 13 + 17 = 50
1 + 3 + 5 + 7 + 15 + 19 = 50
1 + 3 + 5 + 9 + 13 + 19 = 50
1 + 3 + 5 + 9 + 15 + 17 = 50
1 + 3 + 5 + 11 + 13 + 17 = 50
1 + 3 + 7 + 9 + 11 + 19 = 50
1 + 3 + 7 + 9 + 13 + 17 = 50
1 + 3 + 7 + 11 + 13 + 15 = 50
1 + 5 + 7 + 9 + 11 + 17 = 50
1 + 5 + 7 + 9 + 13 + 15 = 50
3 + 5 + 7 + 9 + 11 + 15 = 50

Edwin